Note: For all problems, round percentages to 2 decimal places and currency to the nearest dollar.

PLEASE SHOW ALL STEPS AND HOW YOU OBTAIN THE NUMBERS FOR THE NET PRESENT VALUE FOR EACH PIECE OF EQUIPMENT TO BE PURCHASED. Seeing the calculations will help me have a better understanding. I did not understand this portion of the problem when I first submitted it, so I would appreciate your showing detailed steps on how to solve this portion of this problem. I need this informaton so I can be prepared to complete my pre-exam final. Thank you for your help.

Waters Company is trying to decide which of the two new product lines to introduce in the coming year. The company requires a 10% return on investment. The predicted revenue and cost data for each product line follows:

Product A

Product B

Unit Sales

25,000

20,000

Unit Sales Price

$   35

      $   35

Direct materials

                $    45,000

$       28,000

Direct labor

$ 125,000

$       85,000

Other Cash operating expenses

$    40,000

$       30,000

New equipment costs

$ 2,300,000

$ 1,800,000

Estimated useful life (no salvage)

5 years

5 years

Straight-line depreciation

(2,300,000 – 0)/5 = 460,000

(1,800,000 – 0)/5 =360,000

The company has a 30% tax rate and it used the straight-line depreciation method with no salvage value. The present value of an annuity of $1 for 5 years at 10% is 3.7908.

e. Calculate the net present value for each piece of equipment to be purchased.

                                    Problem A                Problem B
Answer #1
Calculation of expected net cash flow
Particulars Product A Product B
Total sale (A) 875000 700000
less: expenses
Direct material 45000 28000
Direct labor 125000 85000
Operating expenses 40000 30000
Depreciation 460000 360000
Total expenses (B) 670000 503000
Net income before Tax (A)-(B) 205000 197000
Less: tax @ 30% 61500 59100
Net income after tax 143500 137900
add: Depreciation 460000 360000
Expected cash flow 603500 497900
Total sale = sale unit*selling price p.u
Total sale (25000*35),(20000*35) 875000 700000
Calculation of Net Present value
PRODUCT A
Net present value = total present value of expected cash flow-initial investment
Total present value of expected cash flow = expected cash flow*present value annuity of 10%
total present value of expected cash flow = 603500*3.7908 = 2287748
Net present value= 2287748-2300000
Net present value Of product A = -12252.2
PRODUCT B NET PRESENT VALUE
Net present value = total present value of expected cash flow-initial investment
Total present value of expected cash flow = 497900*3.7908
Total present value of expected cash flow = 1887439
NET PRESENT VALUE OF PRODUCT B = 1887439-1800000
NET PRESENT VALUE OF PRODUCT B = 87439.32